Status Report

SOHO Daily Meeting Minutes for Monday, 11 March 2002

By SpaceRef Editor
March 11, 2002
Filed under , ,

* HESSI target for today: AR 9866

* TDRS launched successfully on March 8 so the next SOHO
maneuvers are planned for May 21, allowing for 60 days
of MDI continuous.

Spacecraft Status: Nominal

Spacecraft Anomalies:
069/1052 CDS mnemonic, CUVMCP, violated RH limits for
1 minute 45 seconds. Impacts TBD.

Accomplished Activities:
DOY 068: VIRGO, Flight S/W Maintenance
DOY 069: VIRGO, Noncoho Load

Planned Activities: VIRGO

Upcoming Operations: None

Ground Anomalies:
067/1607 Command corrupted between DSN site and the
spacecraft delayed NRT activities 8 minutes.
070/0806 D46 had to stow antenna due to fire alarm at the
site. Spacecraft placed in record using emergency
record script. Lost 13 min 11 sec of data
recoverable from the SSR.

Other: TDRS successfully launched Friday evening.
MDI continuous resumes on March 19 (first full day).

03/08 DOY 067 – 03/11 DOY 070

EIT daily On-Line Solar Status and observations:

LASCO daily On-Line Solar Status and observations:

Web Page for Planning:

CDS: Nominal. JOP118 on AR 9856 and full sun spectral study today.
Tuesday – JOP118 on AR 9866. An ICAL01 with EIT is planned
for 19 UT Wednesday.

UVCS: Nominal. Coronal hole observations in the south west and long
synoptic. Tuesday – Coronal hole observations at the north
pole and mini synoptic.

LASCO: Nominal. C2/C3 synoptics.

EIT: Nominal. 195 Angstrom CME watch and synoptics.

MDI: Nominal. Full disk magnetograms and dopplergrams.

TRACE: Nominal. Today – Engineering and tracking AR 9859.

SpaceRef staff editor.